The sum of deviations of a set of value x1,x2,x3...xn measured from 50 is -10 and the sum of deviation of the value from 46 is 70 . find the value of n and mean

Grade:10

1 Answers

Zeeshan
6 years ago
In the given set of value there is no class interval Mean=A+ €di/NWhen A= 50, then€di= -10Therefore,Mean = A +€di/N= 50+(-10)/N..........(1)When A= 46,then €di= 70ThereforeAlso mean = A+ €di/N=46+70/N.........(2)From (1) &(2)50 - 10/N=46+70/N4= 80/NN=20Put N= 20 in equity (1), we getMean = 50+ (-10)/20= 49.5 .